# The Symmetry Taco: Equivalences between Gapped, Gapless, and Mixed-State SPTs

### v1

@@ IV.2 Folded Boundaries of Quantum Doubles
Following the preceding general characterization of folded boundaries in $2+1$d TOs described by some UMTC $\mathcal{C}$, let us now specialize to the case when $\mathcal{C}=\mathcal{D}(G)$. We will now translate the constraints that determine when a gapped boundary is an allowed folded boundary into constraints on the $1+1$d gapped phases that can be realized on the folded boundary. Stated physically, we wish to translate these constraints into conditions on the subgroup $H\subset G\times G$ and 2-cocycle class $[\omega]\in H^{2}(H,U(1))$ (which characterize general gapped boundaries of $\mathcal{D}(G\times G)$, see Sec. II) which are compatible with the boundary being a folded boundary of the SymTaco. These conditions will in turn restrict the gapped $1+1$d $G\times G$-symmetric phases which can be realized through the SymTaco  construction and will also allow us to make clear the correspondence between this class of $G\times G$-symmetric gapped phases and $G$-symmetric igSPTS.
@@
As before, we start with the simpler case when $G$ is Abelian. In this case, after folding, we obtain a $\mathcal{D}(G\times G)\cong\mathcal{D}(G)\times\mkern 1.5mu\overline{\mkern-1.5mu\mathcal{D}(G)\mkern-1.5mu}\mkern 1.5mu$ TO with a gapped boundary to vacuum. We can then use the general discussion in the previous subsection to find the conditions under which this gapped boundary is a consistent folded boundary. First, we need to constrain the subgroup $H\subset G\times G$ which, physically, consists of the different flux labels for every anyon that is condensed at the boundary. From the second constraint of Eq. (36), we see that if an anyon $(a,b)\in\mathcal{D}(G)\times\mkern 1.5mu\overline{\mkern-1.5mu\mathcal{D}(G)\mkern-1.5mu}\mkern 1.5mu$ carrying flux $(g_{1},g_{2})\in H$ is condensed, then the anyon $(b,a)$ must also be condensed, implying that $(g_{2},g_{1})$ is also in $H$. Therefore, if the gapped boundary specified by $H$ defines a consistent folded boundary, we can write $H=K\times L$, where $K$ consists of elements of the form $(k,k)$, $L$ consists of elements of the form $(l,e)$, and $L\subset K$. Note that if the condensable subalgebra $\mathcal{A}$ (which is condensed in the middle subregion before folding) contains no pure charges, then $K=G$ [80, 85]. Furthermore, if the condensable algebra $\mathcal{A}$ contains anyons that carry flux in a normal subgroup $A\subset G$, then $L=A$.
@@
Next, we constrain the two-cocycle $\omega$. Recall that, when $H$ is Abelian, the cohomology class of $\omega$ is completely fixed by the function $\Omega(g,h)=\omega(h,g)/\omega(g,h)$ defined in Eq. (3), satisfying the properties listed in Eq. (4). Moreover, as discussed after Eq. (4), the set of anyons in the Lagrangian subgroup specified by $H$ and $\omega$ take the form $(h,\pi_{h})$, where $h\in H$ and $\pi_{h}$ is a $1$d representation of $H$ determined by the function $\Omega(h,\cdot)$. Using the constraints on anyon condensation arising from folded boundaries derived in the previous subsection, we now derive a corresponding constraint on the form of the function $\Omega$.
@@
Therefore, $G\times G$-symmetric folded boundaries of $\mathcal{D}(G\times G)$ can be specified by the functions $\epsilon$ and $\eta$. These functions satisfy the following relations: (a) $\epsilon$ is bilinear, (b) $\eta$ is a 2-cocycle on $L$, and (c) $\epsilon(l,l^{\prime})=\eta(l,l^{\prime})/\eta(l^{\prime},l)$ for all $l,l^{\prime}\in L$. A normal subgroup $L$ and a pair of functions $\eta$, $\epsilon$ satisfying these conditions are exactly the data that specify a condensable subalgebra $\mathcal{A}$ which does not contain any pure charges, as described in Ref. [77] (specializing to the case of Abelian $G$). We see that this same data can be packaged cleanly into a subgroup $H\subset G\times G$ and a 2-cocycle class $[\omega]\in H^{2}(H,U(1))$, which together specify a gapped $1+1$d phase with $G\times G$ symmetry. Thus, we have shown that folded boundaries of the SymTaco  are in one-to-one correspondence with gapped $G\times G$-symmetric $1+1$d phases which are specified by the subgroup $H$ and the 2-cocycle class $[\omega]$ satisfying the constraints derived above.
The above discussion unfortunately does not generalize straightforwardly to the non-Abelian setting. The difficulty primarily arises from the fact that the full data characterizing a gapped boundary cannot be reconstructed given only the set of condensed anyons on the boundary (along with their multiplicities) [83]. The subgroup $H\subset G\times G$ is obtained in the same manner as in the Abelian case discussed above, and consists of elements of the form $(lg,g)$ for $l\in A$ and $g\in G$. This is the subgroup $A\rtimes G$, where the action of $G$ on $A$ is given by conjugation, $l\mapsto glg^{-1}\in A$. However, the 2-cocycle class $[\omega]$ cannot in general be obtained given only the set of condensed anyons and their multiplicities. One must instead resort to the full center construction introduced in Ref. [99] to obtain the data characterizing a folded boundary in the non-Abelian case.
Instead of taking this formal mathematical approach to deriving the fractionalization data for a folded SPT from the corresponding folded boundary, we will utilize a different correspondence made possible by the SymTaco: first, in the following Section, we will use the SymTaco  construction to argue that folded boundaries of $\mathcal{D}(G\times G)$ TOs correspond to $1+1$d $G$-symmetric mixed-state (or average) SPT phases–we will argue this explicitly for $G$ Abelian and conjecture it also holds for $G$ non-Abelian. Next, we will analyze the symmetry fractionalization data characterizing such mixed-state SPTs and use the aforementioned correspondence to obtain the 2-cocycle class $[\omega]$. The result, whose derivation we delay to the next Section, is that
@@
for a folded boundary of $\mathcal{D}(G\times G)$ arising from partial condensation of $\mathcal{A}\subset\mathcal{D}(G)$ described by $\epsilon$ and $\eta$. Under the correspondence to be established between mixed-state SPTs and folded SPTs, this implies that a folded SPT corresponding to the same anyon condensate determined by $(\eta,\epsilon)$ is characterized by the 2-cocycle given by Eq. (44). This (indirectly) implies that folded SPTs have fractionalization data determined by a folded boundary of the corresponding $\mathcal{D}(G\times G)$ SymTaco. Remarkably, this data is identical to the data characterizing $1+1$d $G$-symmetric igSPTs (see Sec. III), which underlines the utility of the SymTaco  construction in obtaining non-trivial correspondences between gapped and gapless phases (albeit with different symmetries), and their corresponding fractionalization data.
In fact, the correspondence established by the SymTaco  between folded SPTs and igSPTs goes beyond a formal identification of the symmetry fractionalization data and provides a direct method for generating folded SPTs from bilayers of igSPTs via suitable Hamiltonian deformations which can be inferred from the SymTaco  construction (as we will discuss in Sec. VI.3, we can use a similar procedure for generating mixed-state SPTs from igSPTs via decoherence that is specified by the SymTaco  construction). Suppose we consider a $G$-symmetric igSPT and stack it with its inverse, i.e., with a $G$-symmetric igSPT which carries the opposite $H^{3}(G_{low},U(1))$ anomaly class (where $G_{low}=G/A$ is the low-energy symmetry group and $A\lhd G$ is a normal subgroup, see Eq. (18)). This stack realizes a bilayer $G\times G$-symmetric $1+1$d system, whose corresponding SymTFT is simply a decoupled stack with $\mathcal{D}(G)\times\mathcal{D}(G)$ on two independent layers. We can then realize a SymTaco  from this bilayer SymTFT by allowing the tunneling of anyons between the two layers, thereby allowing us to gap out the gapless degrees of freedom that remain in the decoupled igSPTs. This corresponds to explicitly breaking the $G\times G$ symmetry down to $A\rtimes G$, consisting of elements of the form $(lg,g)$ for $l\in A$ and $g\in G$. In particular, as discussed in Section II.2, given a reference boundary, the set of anyons condensed on the dynamical boundary of the SymTFT correspond to specific operators in the $1+1$d lattice model; thus, given a Hamiltonian realizing the decopuled igSPT bilayer, the SymTaco, through the diagonal interlayer anyon condensation, suggests the appropriate interlayer couplings to add to the bilayer igSPT Hamiltonian to realize the corresponding folded SPT. We expect that this construction can always be done explicitly on the lattice when $G$ is Abelian. In the following subsection, we illustrate this protocol through the simple example of the $\mathbb{Z}_{4}$ igSPT.
@@ IV.3 Example: $\mathbb{Z}_{4}\times\mathbb{Z}_{4}$ Folded SPT
Recall that the $\mathbb{Z}_{4}$ igSPT can be understood within the SymTFT construction via the condensation of the subalgebra generated by the $e^{2}m^{2}$ boson in the bulk anyon theory $\mathcal{D}(\mathbb{Z}_{4})$, with the fractionalization data provided by Eq. (26). Note that, in contrast to the rest of this section, we label anyons such that an anyon in the second layer, $(1,a)$, has the same spin as the anyon with the same label in the first layer, $(a,1)$, for simplicity. [^11] In other words, here we label anyons $(a,b)\in\mathcal{D}(G)\times\mathcal{D}(G)$ rather than $\mathcal{D}(G)\times\overline{\mathcal{D}(G)}$.. Let us now obtain the folded SPT that corresponds to this igSPT. In this case, we have the SymTaco  defined by the $\mathcal{D}(\mathbb{Z}_{4}\times\mathbb{Z}_{4})$ TO with a gapped boundary to vacuum and we would like to determine the Lagrangian subgroup that corresponds to this gapped boundary (note that we are given the condensable subalgebra $\mathcal{A}$ that determines the intermediate TO as part of the data of the SymTaco).
@@
Here, we used the fact that $[\eta]\in H^{2}(\mathbb{Z}_{2},U(1))=\mathbb{Z}_{1}$ is necessarily trivial, and so we can set $\eta(l,l^{\prime})\equiv 1$. We can now reinterpret this data in terms of the condensed anyons at the folded boundary of the SymTaco  as follows: since the symmetry is broken to the subgroup $H$, the electric anyons $e_{A}^{2}e_{B}^{2}$ are condensed. To determine the charges bound to fluxes, it suffices to consider $\Omega((gl,g),\cdot)$ for the generators of $H$. For the diagonal subgroup generated by $(1,1)$, we have that
@@
Therefore, $\Omega_{(1,1)}$ sends $(g^{\prime},g^{\prime})\mapsto 1$ and $(l^{\prime},1)\mapsto(-1)^{l^{\prime}/2}$. This corresponds to the representation $e_{A}e_{B}^{3}$, which implies that the anyon $e_{A}e_{B}^{3}m_{A}m_{B}$ is condensed. Similarly,
@@
corresponds to the condensation of $e_{A}^{2}$. Therefore, the condensed anyons at the folded boundary are precisely generated by the Lagrangian subgroup in Eq. (15) and the folded SPT in this case corresponds to the third example we had previously encountered in Sec. II.2. Thus, we see that the SymTaco  provides an exact correspondence between the symmetry fractionalization data of this gapped, folded SPT with that of the $1+1$d $\mathbb{Z}_{4}$-symmetric igSPT.

### v2

@@ IV.2 Folded Boundaries of Quantum Doubles
Following the preceding general characterization of folded boundaries in 2+12+1d TOs described by some UMTC $\mathcal{C}$, let us now specialize to the case when $\mathcal{C}=\mathcal{D}(G)$. We will now translate the constraints that determine when a gapped boundary is an allowed folded boundary into constraints on the 1+11+1d gapped phases that can be realized on the folded boundary. Stated physically, we wish to translate these constraints into conditions on the subgroup $H\subset G\times G$ and 2-cocycle class $\omega]\in H^{2}(H,U(1))$ (which characterize general gapped boundaries of $\mathcal{D}(G\times G)$, see Sec. II) which are compatible with the boundary being a folded boundary of the SymTaco. These conditions will in turn restrict the gapped 1+11+1d $G\times G$-symmetric phases which can be realized through the SymTaco  construction and will also allow us to make clear the correspondence between this class of $G\times G$-symmetric gapped phases and $G$-symmetric igSPTS.
@@
As before, we start with the simpler case when $G$ is Abelian. In this case, after folding, we obtain a $\mathcal{D}(G\times G)\cong\mathcal{D}(G)\times\mkern 1.5mu\overline{\mkern-1.5mu\mathcal{D}(G)\mkern-1.5mu}\mkern 1.5mu$ TO with a gapped boundary to vacuum. We can then use the general discussion in the previous subsection to find the conditions under which this gapped boundary is a consistent folded boundary. First, we need to constrain the subgroup $H\subset G\times G$ which, physically, consists of the different flux labels for every anyon that is condensed at the boundary. From the second constraint of Eq. (36), we see that if an anyon $a_{+}b_{-}\in\mathcal{D}(G)\times\mkern 1.5mu\overline{\mkern-1.5mu\mathcal{D}(G)\mkern-1.5mu}\mkern 1.5mu$ carrying flux $(g_{1},g_{2})\in H$ is condensed, then the anyon $b_{+}a_{-}$ must also be condensed, implying that $(g_{2},g_{1})$ is also in $H$. Therefore, if the gapped boundary specified by $H$ defines a consistent folded boundary, we can write $H=K\times L$, where $K$ consists of elements of the form $(k,k)$, $L$ consists of elements of the form $(l,e)$, and $L\subset K$. Note that if the condensable subalgebra $\mathcal{A}$ (which is condensed in the middle subregion before folding) contains no pure charges, then $K=G$ [85, 90]. Furthermore, if the condensable algebra $\mathcal{A}$ contains anyons that carry flux in a normal subgroup $A\subset G$, then $L=A$.
@@
Next, we constrain the two-cocycle $\omega$. Recall that, when $H$ is Abelian, the cohomology class of $\omega$ is completely fixed by the function $\Omega(g,h)=\omega(h,g)/\omega(g,h)$ defined in Eq. (3), satisfying the properties listed in Eq. (4). Moreover, as discussed after Eq. (4), the set of anyons in the Lagrangian subgroup specified by $H$ and $\omega$ take the form $(h,\pi_{h})$, where $h\in H$ and $\pi_{h}$ is a 11d representation of $H$ determined by the function $\Omega(h,\cdot)$. Using the constraints on anyon condensation arising from folded boundaries derived in the previous subsection, we now derive a corresponding constraint on the form of the function $\Omega$.
@@
Therefore, $G\times G$-symmetric folded boundaries of $\mathcal{D}(G\times G)$ can be specified by the functions $\epsilon$ and $\eta$. These functions satisfy the following relations: (a) $\epsilon$ is bilinear, (b) $\eta$ is a 2-cocycle on $L$, and (c) $\epsilon(l,l^{\prime})=\eta(l,l^{\prime})/\eta(l^{\prime},l)$ for all $l,l^{\prime}\in L$. A normal subgroup $L$ and a pair of functions $\eta$, $\epsilon$ satisfying these conditions are exactly the data that specify a condensable subalgebra $\mathcal{A}$ which does not contain any pure charges, as described in Ref. [82] (specializing to the case of Abelian $G$). We see that this same data can be packaged cleanly into a subgroup $H\subset G\times G$ and a 2-cocycle class $\omega]\in H^{2}(H,U(1))$, which together specify a gapped 1+11+1d phase with $G\times G$ symmetry. Thus, we have shown that folded boundaries of the SymTaco  are in one-to-one correspondence with gapped $G\times G$-symmetric 1+11+1d phases which are specified by the subgroup $H$ and the 2-cocycle class $\omega]$ satisfying the constraints derived above.
The above discussion unfortunately does not generalize straightforwardly to the non-Abelian setting. The difficulty primarily arises from the fact that the full data characterizing a gapped boundary cannot be reconstructed given only the set of condensed anyons on the boundary (along with their multiplicities) [88]. The subgroup $H\subset G\times G$ is obtained in the same manner as in the Abelian case discussed above, and consists of elements of the form $(lg,g)$ for $l\in A$ and $g\in G$. This is the subgroup $A\rtimes G$, where the action of $G$ on $A$ is given by conjugation, $l\mapsto glg^{-1}\in A$. However, the 2-cocycle class $\omega]$ cannot in general be obtained given only the set of condensed anyons and their multiplicities. One must instead resort to the full center construction introduced in Ref. [103] to obtain the data characterizing a folded boundary in the non-Abelian case.
Instead of taking this formal mathematical approach to deriving the fractionalization data for a folded SPT from the corresponding folded boundary, we will utilize a different correspondence made possible by the SymTaco: first, in the following Section, we will use the SymTaco  construction to argue that folded boundaries of $\mathcal{D}(G\times G)$ TOs correspond to 1+11+1d $G$-symmetric mixed-state (or average) SPT phases–we will argue this explicitly for $G$ Abelian and conjecture it also holds for $G$ non-Abelian. Next, we will analyze the symmetry fractionalization data characterizing such mixed-state SPTs and use the aforementioned correspondence to obtain the 2-cocycle class $\omega]$. The result, whose derivation we delay to the next Section, is that
@@
for a folded boundary of $\mathcal{D}(G\times G)$ arising from partial condensation of $\mathcal{A}\subset\mathcal{D}(G)$ described by $\epsilon$ and $\eta$. Under the correspondence to be established between mixed-state SPTs and folded SPTs, this implies that a folded SPT corresponding to the same anyon condensate determined by $(\eta,\epsilon)$ is characterized by the 2-cocycle given by Eq. (44). This (indirectly) implies that folded SPTs have fractionalization data determined by a folded boundary of the corresponding $\mathcal{D}(G\times G)$ SymTaco. Remarkably, this data is identical to the data characterizing 1+11+1d $G$-symmetric igSPTs (see Sec. III), which underlines the utility of the SymTaco  construction in obtaining non-trivial correspondences between gapped and gapless phases (albeit with different symmetries), and their corresponding fractionalization data.
In fact, the correspondence established by the SymTaco  between folded SPTs and igSPTs goes beyond a formal identification of the symmetry fractionalization data and provides a direct method for generating folded SPTs from bilayers of igSPTs via suitable Hamiltonian deformations which can be inferred from the SymTaco  construction (as we will discuss in Sec. VI.3, we can use a similar procedure for generating mixed-state SPTs from igSPTs via decoherence that is specified by the SymTaco  construction). Suppose we consider a $G$-symmetric igSPT and stack it with its inverse, i.e., with a $G$-symmetric igSPT which carries the opposite $H^{3}(G_{low},U(1))$ anomaly class (where $G_{low}=G/A$ is the low-energy symmetry group and $A\lhd G$ is a normal subgroup, see Eq. (18)). This stack realizes a bilayer $G\times G$-symmetric 1+11+1d system, whose corresponding SymTFT is simply a decoupled stack with $\mathcal{D}(G)\times\mathcal{D}(G)$ on two independent layers. We can then realize a SymTaco  from this bilayer SymTFT by allowing the tunneling of anyons between the two layers, thereby allowing us to gap out the gapless degrees of freedom that remain in the decoupled igSPTs. This corresponds to explicitly breaking the $G\times G$ symmetry down to $A\rtimes G$, consisting of elements of the form $(lg,g)$ for $l\in A$ and $g\in G$. In particular, as discussed in Section II.2, given a reference boundary, the set of anyons condensed on the dynamical boundary of the SymTFT correspond to specific operators in the 1+11+1d lattice model; thus, given a Hamiltonian realizing the decopuled igSPT bilayer, the SymTaco, through the diagonal interlayer anyon condensation, suggests the appropriate interlayer couplings to add to the bilayer igSPT Hamiltonian to realize the corresponding folded SPT. We expect that this construction can always be done explicitly on the lattice when $G$ is Abelian. In the following subsection, we illustrate this protocol through the simple example of the $\mathbb{Z}_{4}$ igSPT.
@@ IV.3 Example: $\mathbb{Z}_{4}\times\mathbb{Z}_{4}$ Folded SPT
Recall that the $\mathbb{Z}_{4}$ igSPT can be understood within the SymTFT construction via the condensation of the subalgebra generated by the $e^{2}m^{2}$ boson in the bulk anyon theory $\mathcal{D}(\mathbb{Z}_{4})$, with the fractionalization data provided by Eq. (26). Here, we label anyons $a_{A}b_{B}\in\mathcal{D}(G)\times\mathcal{D}(G)$ rather than $\mathcal{D}(G)\times\overline{\mathcal{D}(G)}$. Let us now obtain the folded SPT that corresponds to this igSPT. In this case, we have the SymTaco  defined by the $\mathcal{D}(\mathbb{Z}_{4}\times\mathbb{Z}_{4})$ TO with a gapped boundary to vacuum and we would like to determine the Lagrangian subgroup that corresponds to this gapped boundary (note that we are given the condensable subalgebra $\mathcal{A}$ that determines the intermediate TO as part of the data of the SymTaco).
@@
Here, we used the fact that $\eta]\in H^{2}(\mathbb{Z}_{2},U(1))=\mathbb{Z}_{1}$ is necessarily trivial, and so we can set $\eta(l,l^{\prime})\equiv 1$. We can now reinterpret this data in terms of the condensed anyons at the folded boundary of the SymTaco  as follows: since the symmetry is broken to the subgroup $H$, the electric anyons $e_{A}^{2}e_{B}^{2}$ are condensed. To determine the charges bound to fluxes, it suffices to consider $\Omega((gl,g),\cdot)$ for the generators of $H$. For the diagonal subgroup generated by $(1,1)$, we have that
@@
Therefore, $\Omega_{(1,1)}$ sends $(g^{\prime},g^{\prime})\mapsto 1$ and $(l^{\prime},1)\mapsto(-1)^{l^{\prime}/2}$. This corresponds to the representation $e_{A}e_{B}^{3}$, which implies that the anyon $e_{A}m_{A}e_{B}^{3}m_{B}$ is condensed. [^11] Note that this is the image of $e_{+}m_{+}\bar{e}_{-}\bar{m}_{-}$ under the isomorphism $\overline{\mathcal{D}(G)}\to\mathcal{D}(G\times G)$. Similarly,
@@
corresponds to the condensation of $e_{A}^{2}$. Therefore, the condensed anyons at the folded boundary are precisely generated by the Lagrangian subgroup in Eq. (15) and the folded SPT in this case corresponds to the third example we had previously encountered in Sec. II.2. Thus, we see that the SymTaco  provides an exact correspondence between the symmetry fractionalization data of this gapped, folded SPT with that of the 1+11+1d $\mathbb{Z}_{4}$-symmetric igSPT.
@@
$+$ $\mathcal{Z}_{A,i-1}\mathcal{Z}^{\dagger}_{A,i}\mathcal{X}_{A,i}\mathcal{Z}_{B,i-1}\mathcal{Z}^{\dagger}_{B,i}\mathcal{X}_{B,i}+\text{ h.c.}\Big)\,.$
The first two terms impose the energetic constraints Eq. (32) on the two copies of the system and correspond to the condensation of $e^{2}m^{2}$ and $\bar{e}^{2}\bar{m}^{2}$ on each copy in the corresponding $\mathcal{D}(\mathbb{Z}_{4})$ and $\mkern 1.5mu\overline{\mkern-1.5mu\mathcal{D}(\mathbb{Z}_{4})\mkern-1.5mu}\mkern 1.5mu$ SymTFTs. The second line constructs the SymTaco  from the decoupled layers and corresponds to tunneling $e^{2}$ between the two copies (equivalently, condensing $e^{2}\bar{e}^{2}$), and spontaneously breaks the original $\mathbb{Z}_{4}\times\mathbb{Z}_{4}$ symmetry to $\mathbb{Z}_{2}^{A}\times\mathbb{Z}_{4}^{AB}$. The final term fully gaps out the system and corresponds to tunneling $em$ and $e^{3}m$ between the two copies. The ground of this Hamiltonian realizes the non-trivial 1+11+1d gapped $\mathbb{Z}_{4}\times\mathbb{Z}_{2}$ SPT that was discussed as Example 3 in Sec. II.2.
V Mixed-State SymTFT: Folded Boundaries as Mixed-State SPTs